balance the following redox reactions and indentify the reducing agent and oxidizing agent
Cd2+(aq) + Cu+(aq) arrow Cd(s) + Cu+2(aq)
Al3+(aq) + Co(s) arrow Al(s) + Co2+(aq)
1
Expert's answer
2019-11-21T06:25:35-0500
  1. Cd(s)+ 2cu+(aq) =Cd2+(aq) +2cu(s) .In this case,Cu is the oxidizing agent and Cd is the reducing agent.
  2. Al3+(aq) +2Co(s) =3Al(s) + Co2+(aq). Co is the reducing agent whereas Al is the oxidizing agent.